OpenStudy (anonymous):

One student can proofread a copy for the newspaper in 4 hours. if a second proofreader is also employed, the job can be done in 2.5 hours. how long does it take for the second proofreader to do the same job alone?

OpenStudy (kropot72):

The student who proofreads a copy in 4 hours can read 1/4 = 0.25 of the copy in 1 hour. When two people are proofreading, 1/2.5 = 0.4 of the copy is read in one hour. Therefore the second proofreader reads 0.4 - 0.25 = 0.15 of the copy in 1 hour. Now we can find the time for the second proofreader to do the same job alone by taking the reciprocal of 0.15: \[\frac{1}{0.15}=?\ hours+minutes\]

OpenStudy (kropot72):

@lior94 Are you there? Do you understand?

OpenStudy (anonymous):

yes, im here. i mot sure how you got .25 for the first reader though, could you explain please?

OpenStudy (kropot72):

If the first reader takes 4 hours to read the copy, the reader will read a quarter of the copy in one hour. 1/4 = 0.25